# MNO: Multiscale Neural Operator for Computational Fluid Dynamics with 3D Point Cloud Data

### v1

@@ 3.4 Micro Point-wise Attention Module
Micro-scale features correspond to the intrinsic attributes of individual spatial points. To complement the global and local attention modules, which capture long-range and neighborhood-level interactions, the micro point-wise attention module focuses on the single-point scale that could be overlooked at coarser scales.
This module implements a point-wise self-attention mechanism, where each token is reweighted solely based on its own feature vector. As illustrated in Figure 1(e), token features from the previous block are processed through an MLP followed by a softmax operation to produce point-specific weights, which indicate the relative importance of each token. The scaled features are then combined with the original input via a residual connection
$\begin{split}X_{micro}=X+Score_{p}\odot X,\;\text{ with }\;Score_{p}=\text{Softmax}(\text{MLP}(X),\text{ dim}=N),\end{split}$ (5)
where $X\in\mathbb{R}^{N\times D}$ denotes the input token features, $Score_{p}\in\mathbb{R}^{N\times 1}$ represents the point-wise attention weights, and $X_{micro}\in\mathbb{R}^{N\times D}$ is the resulting micro-scale representation. The symbol $\odot$ indicates element-wise multiplication with broadcasting rule across feature dimensions.
Because each token is reweighted independently of others, this module emphasizes point-specific high-frequency variations while remaining computationally efficient, with complexity $O(ND^{2})$. In combination with the global and local attentions, it ensures that MNO simultaneously preserves fine-grained details, local neighborhood patterns, and long-range flow coherence.
Remarks on other multiscale models
Existing multiscale methods for point clouds primarily focus on visual tasks, e.g., segmentation (Qi et al., 2017b; Hu et al., 2020; Yu et al., 2018). These approaches typically rely on repeated downsampling and upsampling, which can discard fine-grained information and lead to suboptimal accuracy when applied to flow field prediction (Qi et al., 2017b).
@@ 4.1 Benchmarks
We evaluate the model performance on four 3D CFD benchmarks, including steady-state flow field benchmarks, Ahmed body (Ahmed et al., 1984; Li et al., 2023b), ShapeNet Car (Umetani & Bickel, 2018), DrivAerNet++ (Elrefaie et al., 2024), and the unsteady flow field benchmark, Parachute dynamics (Zeng et al., 2025).
Ahmed body (100k/sample): A vehicle wind tunnel dataset with a bluff-body structure. Inputs consist of the vehicle surface point cloud and auxiliary conditions such as freestream velocity and Reynolds number. The output is the pressure field on the vehicle surface. Parachute dynamics (15k/sample): A time-dependent dataset capturing the inflation of parachutes under pressure loads. Inputs include the initial point cloud positions and markers for the umbrella surface and ropes, while outputs are displacement fields at four time steps. ShapeNet Car (30k/sample): A car wind tunnel dataset. Inputs include point positions, signed distance values, and surface normals. Outputs are the velocity field in the air region and the pressure field on the car surface. DrivAerNet++ (300k/sample): A large-scale automotive wind tunnel dataset. Inputs consist of point positions and surface normals, and the output is the pressure field on the car surface. For detailed configurations, please refer to Appendix B.
4.2 Comparison on Accuray
We reproduce several state-of-the-art open-source methods of neural operator for comparative experiments. To ensure fairness, all comparative experiments are conducted on the same protocol and executed on our identical computational hardware.
Table 1: The comparison results with other methods on Ahmed body and Parachute datasets, in which $RL2_{p}$ denotes the relative $L_{2}$ errors (RL2) of the pressure field; $RL2_{x1\sim 4}$ represent the RL2 of the displacement field at 4 time steps; $RL2_{x}$ denotes the total RL2 of 4 time steps; MAEMAE is the mean absolute errors. The subscript “*” indicates the result claimed in the original article. The row titled with Improvement refers to the degree of advancement compared to the previous best method.
@@ 4.2 Comparison on Accuray
LNO (Wang & Wang, 2024) 0.1908 30.4570 0.0584 0.0431 0.0484 0.0665 0.0504 0.0147
PCNO (Zeng et al., 2025) 0.0682 - - - - - 0.0373 -
PCNO (Zeng et al., 2025) 0.0664 12.4693 0.0238 0.0189 0.0305 0.0515 0.0316 0.0094
@@
Table 2: The comparison results with other advanced methods on ShapeNet Car and DrivAerNet++ datasets. $RL2_{v}$ denotes the RL2 of the velocity field. The subscript “*” indicates the result claimed in the original article.
ShapeNet Car DrivAerNet++
Methods $RL2_{p}$ $MAE_{p}$ $RL2_{v}$ $MAE_{v}$ $RL2_{p}$ $MAE_{p}$
DeepONet (Lu et al., 2021) 0.4148 11.4996 0.2075 1.2256 0.3203 27.4931
PointNet (Qi et al., 2017a) 0.0927 2.6222 0.0314 0.1723 0.4278 42.6893
PointNet++ (Qi et al., 2017b) 0.2082 5.9648 0.0771 0.3813 0.4617 41.8497
Geo-FNO (Li et al., 2023a) 0.1164 3.5748 0.0737 0.4647 0.2869 26.3732
LNO (Wang & Wang, 2024) 0.0887 2.6118 0.0267 0.1498 0.1984 18.1088
Transolver (Wu et al., 2024) 0.0745 - 0.0207 - - -
Transolver (Wu et al., 2024) 0.0700 1.8151 0.0230 0.1130 0.1749 15.4372
Ours 0.0597 1.3796 0.0178 0.0845 0.1665 14.6335
Improve 14.71% 23.99% 22.61% 25.22% 4.80% 5.21%
The methods we compared are commented as follows.
DeepONet (Lu et al., 2021): The first neural operator framework, primarily based on MLPs. Its representational capacity is limited, particularly when addressing highly nonlinear 3D PDEs.
PointNet’s (Qi et al., 2017a): Utilizes global pooling to model the overall structure of flow fields. However, it neglects local interactions among points, and in dense point clouds, the pooling mechanism may excessively compress information.
PointNet++ (Qi et al., 2017b): Extends PointNet with fixed-number downsampling to extract multiscale features. On large-scale point clouds, this strategy may discard essential geometric structures, sometimes resulting in performance worse than PointNet. Moreover, increasing the sampled points significantly raises computational cost and parameter count, making the method less practical.
Geo-FNO (Li et al., 2023a): Maps irregular grids onto regular ones and applies FNO for global Fourier-based convolutions. Its performance is highly dependent on the quality of the mapping and lacks robustness for complex geometries.
Transolver (Wu et al., 2024): Employs a global dimension reduction and introduces residual branch between token compression and decompression to reduce information loss. However, it does not explicitly support multiscale feature learning.
LNO (Wang & Wang, 2024): Compresses point clouds into a latent space with limited tokens, where multiple Transformer layers capture global features. This approach resembles our Global Attention module but suffers from noticiable loss of fine-grained details due to heavy compression.
PCNO (Zeng et al., 2025): Extracts global, gradient, and residual features of the input point cloud. However, its global feature extraction relies on FNO without point cloud compression, limiting scalability for large datasets. Compared to PCNO, MNO provides stronger mid-scale representations through its Local Attention module.
@@
The comparative results are summarized in Table 1 and Table 2. The proposed MNO consistently delivers higher predictive accuracy across all four benchmarks compared to recent baselines. In particular, relative to the current leading methods, Transolver and PCNO, MNO achieves error reductions of 29.51% on the Ahmed Body dataset, 15.82% on Parachute Dynamics, 14.71% on ShapeNetCar, and 4.80% on DrivAerNet++.
@@ 4.3 Attention Modules Ablation Experiments
In each MNO block, the three attention modules are responsible for extracting multiscale features. To better understand their contributions, we conduct ablation studies by selectively enabling different modules. Unless otherwise specified, the number of MNO blocks is fixed at four.
The results are summarized in Table 3. “Global,” “Local,” and “Micro” denote using only the corresponding attention module to learn and predict flow fields. “Global+Local” indicates the joint use of both Global and Local Attention modules, “Global+Local+Micro” represents the full MNO block, and “Global+Global+Global” refers to using three identical Global Attention modules as a control.
@@
Ahmed body Parachute ShapeNet Car DrivAerNet++
Modules $RL2_{p}$ $MAE_{p}$ $RL2_{x}$ $MAE_{x}$ $RL2_{p}$ MAE_p $RL2_{v}$ $MAE_{v}$ $RL2_{p}$ $MAE_{p}$
Global 0.8588 154.899 0.8205 0.2629 0.5117 16.9729 0.2025 1.3195 0.7853 73.3534
Local 0.1350 24.8293 0.1479 0.0314 0.0832 2.9235 0.0399 0.2042 0.1919 17.8505
Micro 0.4028 64.5137 0.2307 0.0542 0.1881 6.2470 0.0609 0.3098 0.2396 21.2415
Local+Micro 0.1267 24.0813 0.1062 0.0234 0.0807 2.7964 0.0393 0.1975 0.1908 17.6443
Global+Micro 0.0484 7.4700 0.0304 0.0095 0.0663 1.5408 0.0194 0.0980 0.1728 15.3040
Global+Local 0.0488 7.6412 0.0287 0.0090 0.0610 1.4994 0.0201 0.0983 0.1713 14.9961
Global+Global+Global 0.8591 153.907 0.8170 0.2601 0.7986 23.7095 0.3249 2.0015 0.7853 73.2887
Local+Local+Local OOM OOM 0.1484 0.0313 0.0800 2.7651 0.0385 0.1852 OOM OOM
Micro+Micro+Micro 0.4328 76.6320 0.2249 0.0527 0.1893 6.1662 0.0591 0.2930 0.2420 21.5583
Global+Local+Micro 0.0468 7.0465 0.0266 0.0081 0.0597 1.3796 0.0178 0.0845 0.1665 14.6335
@@
When used individually, the Global Attention module performs worst due to repeated low-rank projections that discard point cloud information. By contrast, the Local Attention module achieves the best performance among the three. Local Attention captures mid-scale features, i.e., mid-frequency information, which is crucial for distinguishing geometric shapes of objects.
@@
Combining Global and Local Attention substantially improves performance, with relative gains of 63.85%, and 80.59%, 25.33%, and 10.73%, across the four datasets compared to using Local Attention alone. This highlights the strong complementarity of global and mid-scale features, showing that their combination captures most of the key physical processes in flow fields.
@@
Adding Micro Attention on top of Global+Local yields further improvements of 4.09%, 7.32%, 2.13%, and 2.80%, across the four datasets. The Micro Attention module captures high-frequency variations that serve as fine-scale corrections to mid-frequency features. While its contribution is smaller, it refines predictions and enhances overall accuracy.
4.4 Visualization of Attention Modules

### v2

@@ 3.4 Micro Point-wise Attention Module
The micro-scale features correspond to the intrinsic attributes of individual spatial points. This module implements a point-wise self-attention mechanism, where each token is reweighted solely based on its own feature vector. As illustrated in Figure 1 (e), token features from the previous block are processed through an MLP followed by a softmax operation to produce point-specific weights, which indicate the relative importance of each token. The scaled features are then combined with the original input via a residual connection
$\begin{split}X_{micro}&=X+Score_{p}\odot X,\ X\in\mathbb{R}^{N\times D},\\ Score_{p}&=\text{Sigmoid}(\text{MLP}(X)),\ Score_{p}\in\mathbb{R}^{N\times D},\end{split}$ (6)
where $X$ denotes the input token features; $Score_{p}$ represents the point-wise attention weights, and $X_{micro}\in\mathbb{R}^{N\times D}$ is the resulting micro-scale representation. The symbol $\odot$ indicates element-wise multiplication with broadcasting rule across feature dimensions.
The Local Module leverages differential attention to model inter-node relational variations (equation (5)), while the Micro Module restores absolute point-wise states that are annihilated by differential operators. Together, the mid- and high-frequency representations learned by the Local and Micro Modules substantially enhance the Global Module’s fine-grained detail perception. The three modules are thus tightly coupled and jointly enable more expressive modeling of physical spatiotemporal flow fields in MNO.
Remarks on other multiscale models: Existing multiscale methods for point clouds primarily focus on multi-level sampling operation (Li et al., 2020, 2025; Qi et al., 2017b; Hu et al., 2020). Repeated downsampling and upsampling discard fine-grained geometric details, yielding suboptimal accuracy in flow field prediction. Moreover, neighborhood relationships must be recomputed after each sampling step, incurring substantial computational overhead. The results in Table 6 further confirm this conclusion.
@@ 4.1 Benchmarks
We evaluate the model performance on five 3D CFD benchmarks, including steady-state flow field benchmarks, Ahmed body (Ahmed et al., 1984; Li et al., 2023b), ShapeNet Car (Umetani and Bickel, 2018), DrivAerNet++ (Elrefaie et al., 2024), and the unsteady flow field benchmark, Parachute dynamics (Zeng et al., 2025b), Magnetohydrodynamics (MHD) (Ohana et al., 2024).
Ahmed body (100k/sample): A vehicle wind tunnel dataset with a bluff-body structure. Inputs consist of the vehicle surface point cloud and auxiliary conditions such as freestream velocity and Reynolds number. The output is the pressure field on the vehicle surface. Parachute dynamics (15k/sample): A time-dependent dataset capturing the inflation of parachutes under pressure loads. Inputs include the initial point cloud positions and markers for the umbrella surface and ropes, while outputs are displacement fields at four time steps. MHD (260k/sample): A 3D magnetohydrodynamic turbulence dataset. The input comprises observed density, velocity, and magnetic fields from four consecutive time steps, with the goal of predicting all fields at the next time step. ShapeNet Car (30k/sample): A car wind tunnel dataset. Inputs include point positions, signed distance values, and surface normals. Outputs are the velocity field in the air region and the pressure field on the car surface. DrivAerNet++ (300k/sample): A large-scale automotive wind tunnel dataset. Inputs consist of point positions and surface normals, and the output is the pressure field on the car surface. For detailed configurations, please refer to Appendix B.
Table 1: The comparison results with other methods on Ahmed body and Parachute datasets, in which $RL2_{p}$ denotes the relative $L_{2}$ errors (RL2) of the pressure field; $RL2_{x1\sim 4}$ represent the RL2 of the displacement field at 4 time steps; $RL2_{x}$ denotes the total RL2 of 4 time steps; MAEMAE is the mean absolute errors. The subscript “*” indicates the result claimed in the original article. The row titled with “Improvement” refers to the degree of advancement compared to the previous best method.
@@
LNO (Wang and Wang, 2024) 0.1908 30.4570 0.0584 0.0431 0.0484 0.0665 0.0504 0.0147
AMG (Li et al., 2025) - - 0.0432 0.0288 0.0369 0.0539 0.0379 0.0120
PCNO (Zeng et al., 2025b) 0.0682 - - - - - 0.0373 -
PCNO (Zeng et al., 2025b) 0.0664 12.4693 0.0238 0.0189 0.0305 0.0515 0.0316 0.0094
@@
Table 2: The comparison results with other advanced methods on Magnetohydrodynamics, ShapeNet Car and DrivAerNet++ datasets. $RL2_{d}$, $RL2_{v}$, $RL2_{m}$, and $RL2$ represent the RL2 of density, velocity, magnetic fields, and their average value.
Magnetohydrodynamics (MHD) ShapeNet Car DrivAerNet++
Methods $RL2_{d}$ $RL2_{v}$ $_{m} RL2RL2$ MAEMAE $RL2_{p}$ $MAE_{p}$ $RL2_{v}$ $MAE_{v}$ $RL2_{p}$ $MAE_{p}$
DeepONet(Lu et al., 2021) 0.4855 1.0004 1.0013 0.8290 0.6193 0.4148 11.4996 0.2075 1.2256 0.3203 27.4931
PointNet(Qi et al., 2017a) 0.2304 0.2771 0.3316 0.2797 0.1855 0.0927 2.6222 0.0314 0.1723 0.4278 42.6893
PointNet++(Qi et al., 2017b) 0.4343 0.6397 0.6647 0.5795 0.4215 0.2082 5.9648 0.0771 0.3813 0.4617 41.8497
Geo-FNO(Li et al., 2023a) 0.5196 1.0043 0.9955 0.8398 0.6255 0.1164 3.5748 0.0737 0.4647 0.2869 26.3732
LNO(Wang and Wang, 2024) 0.2310 0.2675 0.3303 0.2762 0.1826 0.0887 2.6118 0.0267 0.1498 0.1984 18.1088
AMG(Li et al., 2025) - - - - - 0.0770 2.0062 0.0236 0.1203 - -
Transolver(Wu et al., 2024) - - - - - 0.0745 - 0.0207 - - -
Transolver(Wu et al., 2024) 0.2282 0.2655 0.3274 0.2737 0.1804 0.0700 1.8151 0.0230 0.1130 0.1749 15.4372
Ours 0.1031 0.1354 0.1378 0.1254 0.08496 0.0597 1.3796 0.0178 0.0845 0.1665 14.6335
Improve 54.82% 49.00% 57.91% 54.18% 52.90% 14.71% 23.99% 22.61% 25.22% 4.80% 5.21%
Figure 2: Prediction results with identical model parameter counts. (a) the pressure error on ShapeNet Car, (b) the velocity error on ShapeNet Car, (c) the displacement RL2 error on Parachute, and (d) the displacement MAE on Parachute.
4.2 Comparison on Accuracy
We reproduce several state-of-the-art open-source methods of neural operator for comparative experiments. The training and testing procedures for all baselines are consistent with MNO. The hyperparameters of baselines adhere to their official code repositories or original papers. Detailed settings can be found in Appendix C.
@@ 4.2 Comparison on Accuracy
The comparative results are summarized in Table 1 and Table 2. The proposed MNO consistently delivers higher predictive accuracy across all five benchmarks compared to recent baselines. In particular, relative to the current leading methods, Transolver and PCNO, MNO achieves error reductions of 29.51% on the Ahmed Body dataset, 15.82% on Parachute Dynamics, 54.18% on Magnetohydrodynamics, 14.71% on ShapeNetCar, and 4.80% on DrivAerNet++. Notably, on the challenging 3D turbulent spatiotemporal prediction task, MHD, methods based solely on global feature (Geo-FNO, LNO, Transolver) exhibit clear performance degradation, whereas MNO achieves an improvement of about 50%. This can be attributed to the strong nonlinearity and pronounced multi-scale local turbulent structures in MHD, which are difficult to capture with purely global modeling.
Below, we provide an analysis and discussion of some key factors underlying MNO’s methodology over state-of-the-art methods that leading better results. LNO (Wang and Wang, 2024) compresses point clouds into a latent space with limited tokens, where multiple Transformer layers capture global features. This approach resembles our Global Attention module but suffers from noticiable loss of fine-grained details due to heavy compression. Transolver (Wu et al., 2024) employs a global dimension reduction and introduces residual branch between token compression and decompression to reduce information loss. However, it does not explicitly support multiscale feature learning. PCNO (Zeng et al., 2025b) extracts global, gradient, and residual features of the input point cloud. However, its global feature extraction relies on FNO without point cloud compression, limiting scalability for large datasets. Compared to PCNO, MNO provides stronger mid-scale representations through its Local Attention module. AMG (Li et al., 2025) constructs global and local graph structures through multiple downsampling, but this results in significant loss of spatial details, thereby limiting prediction accuracy. Moreover, repeated farthest point sampling (FPS) significantly increases computation time.
4.3 Comparison on Model Scales
This experiment aims to demonstrate that, compared with other baselines, the improvements of our MNO model are primarily due to its innovative design of multiscale structure rather than merely an increase in model scale.
To ensure fairness, all models are compared under consistent parameter scales. For our MNO model, the number of blocks is set to 1, 2, 4, and 8 to construct four parameter scales. For Transolver (Wu et al., 2024), the token length is set to 264, and the network depth is adjusted to 1, 2, 4, and 8 to align with the parameter scales. For LNO (Wang and Wang, 2024), the token length is set to 88, 120, 168, and 240 to match the parameter scales. For AMG (Li et al., 2025), the token length is set to 88, 116, 162 and 228 to match the parameter scales.
The experimental results, as shown in Figure 2, indicate that under the same parameter scales, the prediction error of our MNO model is consistently lower than that of the baselines. For example, in Figure 2 (a), our model achieved prediction errors that are 13%-20% lower than the best baseline across the four parameter scales. When the RL2 error is below 0.075, the parameter scale of our model is reduced by 46.61% compared to the best baseline. This indicates that MNO does not need to design redundant parameter spaces for the underlying physical state of the flow field, as adopted by the baselines in Fig. 2.
@@ 4.4 Attention Modules Ablation Experiments
In each MNO block, the three attention modules are responsible for extracting multiscale features. To better understand their contributions, we conduct ablation studies by selectively enabling different modules.
The results are summarized in Table 3. The ablation visualization of these three attention modules is shown in Appendix A. “Global,” “Local,” and “Micro” denote using only the corresponding attention module to learn and predict flow fields. “Global+Local” indicates the joint use of both Global and Local Attention modules, “Global+Local+Micro” represents the full MNO block. The “Global+Global+Global” refers to using three identical Global Attention modules to show the improvement stems from the complementary benefits of our multi-scale design, rather than any single type of model. Unless otherwise specified, the number of MNO blocks is fixed at four.
@@
Ahmed body Parachute MHD ShapeNet Car DrivAerNet++
Modules $RL2_{p}$ $MAE_{p}$ $RL2_{x}$ $MAE_{x}$ $RL2$ MAEMAE $RL2_{p}$ MAE_pMAE\_p $RL2_{v}$ $MAE_{v}$ $RL2_{p}$ $MAE_{p}$
Global 0.8588 154.899 0.8205 0.2629 0.7597 0.5576 0.5117 16.9729 0.2025 1.3195 0.7853 73.3534
Local 0.1350 24.8293 0.1479 0.0314 0.1712 0.1163 0.0832 2.9235 0.0399 0.2042 0.1919 17.8505
Micro 0.4028 64.5137 0.2307 0.0542 0.2728 0.1797 0.1881 6.2470 0.0609 0.3098 0.2396 21.2415
Local+Micro 0.1267 24.0813 0.1062 0.0234 0.1344 0.0911 0.0807 2.7964 0.0393 0.1975 0.1908 17.6443
Global+Micro 0.0484 7.4700 0.0304 0.0095 0.2722 0.1795 0.0663 1.5408 0.0194 0.0980 0.1728 15.3040
Global+Local 0.0488 7.6412 0.0287 0.0090 0.1529 0.1041 0.0610 1.4994 0.0201 0.0983 0.1713 14.9961
Global+Global+Global 0.8591 153.907 0.8170 0.2601 0.7603 0.5558 0.7986 23.7095 0.3249 2.0015 0.7853 73.2887
Local+Local+Local OOM OOM 0.1484 0.0313 OOM OOM 0.0800 2.7651 0.0385 0.1852 OOM OOM
Micro+Micro+Micro 0.4328 76.6320 0.2249 0.0527 0.2730 0.1799 0.1893 6.1662 0.0591 0.2930 0.2420 21.5583
Global+Local+Micro 0.0468 7.0465 0.0266 0.0081 0.1254 0.08496 0.0597 1.3796 0.0178 0.0845 0.1665 14.6335
@@
Performance of independent modules: Global-only models suffer from severe information loss due to cascaded low-rank projections, resulting in degenerated predictions. By contrast, the Local Attention module achieves the best performance among the three. Local Attention captures mid-scale features, i.e., mid-frequency information, which is crucial for distinguishing geometric shapes of objects.
@@
Improvement of global module: Combining Global and Local Attention substantially improves performance, with relative gains of 63.85%, 80.59%, 26.68%, 10.68%, and 10.73%, across the five datasets compared to using Local Attention alone. This highlights the strong complementarity of global-scale and mid-scale features, showing that their combination captures most of the key physical processes in flow fields.
@@
Improvement of micro module: Adding Micro Attention on top of Global+Local yields further improvements of 4.09%, 7.32%, 17.98%, 2.13%, and 2.80%, across the five datasets. The Micro Attention module captures high-frequency variations that serve as fine-scale corrections to mid-frequency features. While its contribution is smaller, it refines predictions and enhances overall accuracy.
Coupling of multi-scale modules: The performance of individual modules is limited, while their combination can produce powerful results, which is the design advantage of MNO. Local and Micro modules provide detailed local spatial information to supplement global modules, while global modules provide a broader perspective to enhance Local and Micro units. Therefore, any combination of them can improve performance. These three modules work together to improve the overall performance of MNO, without the need for each module to handle CFD tasks independently. Conversely, if the Global and Micro modules performed well independently, it would indicate architectural redundancy rather than effective specialization.
Other experiments: The computing cost is shown in Appendix D. Ablation on the number of MNO blocks, the $M$, and the $k_{nbr}$ are detailed in Appendix E, F and G. The evaluation of the MNO’s adaptive resolution capabilities and geometric generalization are detailed in Appendices H and I. For a comparative visualization of the key local flow fields between MNO and the baseline, refer to Appendix K.
5 Conclusions
Impact Statement
This paper aims to accelerate high fidelity industrial simulations, such as those for automobiles, aircraft, and wind turbines, via multi scale neural operators. Our work may have various potential societal impacts; however, we believe that none require particular emphasis here.
Appendix A Visualization of Attention Modules